Monday, 20 August 2007

Week 33 start of 34

Main goings on have been the DMP update this was hampered by password version problems in MySQL, was resolved with help from here . More problem came with the change to BDii new firewall ports open and yet there was still no data coming out.

The BDii was going to be fixed today, however Sheffield has suffered several power cut over the last 24 hours. This has affected the hole of the LCG here, recovery work is ongoing.

No comments: